# Chain-of-Thought in Neural Code Generation: From and For Lightweight Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable potential in code generation. The integration of Chain of Thought (CoT) reasoning can further boost their performance. However, current CoT methods often require manual writing or LLMs with over 100 billion parameters to generate, impeding their applicability in resource-constrained scenarios. In this study, we investigate lightweight Language Models ($\ell$LMs), which are defined to have fewer than 10 billion parameters. Empirically, we find that most $\ell$LMs cannot generate high-quality CoTs when prompted by the few-shot method, but can take advantage of high-quality CoTs generated elsewhere to improve their performance in code generation. Based on these findings, we design a novel approach COTTON which can leverage $\ell$LMs to automatically generate CoTs for code generation. We synthesize new datasets and conduct extensive experiments on various benchmarks. The results show that the CoTs generated by COTTON outperform the baselines in terms of automated and human evaluation metrics. In particular, the CoTs generated by COTTON boost various $\ell$LMs to achieve higher performance gains than those generated by LLMs such as ChatGLM (130B), and are competitive with those generated by gpt-3.5-turbo (175B). Our study also showcases the potential of $\ell$LMs in software engineering applications.
@@ 1 Introduction
Neural code generation, which can automatically generate programs from natural language requirements based on deep learning has become a promising approach to meet the challenges of the ever-increasing complexity of software and alleviate the burden on programmers [1, 2]. Recently, large language models (LLMs), such as GPT4 [3], have demonstrated impressive performance in code generation tasks [4]. The state-of-the-art LLMs normally have over 100 billion parameters, making even their deployment highly non-trivial. These LLMs pose challenges in terms of time, computational, and financial costs when applied to code generation, rendering them impractical for most individual users, or in resource-constrained scenarios, such as restricted access to LLM APIs or constrained GPU availability. [5, 6]. For software engineering applications, it is imperative to develop lightweight language models based techniques that are more friendly for users (e.g., individual end users). In this study, we refer to the study of Fu et al. [7] and define language models (LMs) with less than 10 billion parameters as lightweight Language Models ($\ell$LM), the rationale of which is that they can be deployed on a single user graphics card (e.g., RTX 3090 or RTX 4090) based on the current technology. The general aim is to develop techniques to tackle software engineering challenges based on $\ell$LMs but with competitive performance as state-of-the-art LLMs, which would enable efficient, yet more accessible, software engineering applications.
@@
An example of the CoT prompt is shown in Figure 1(b), where the programming task is choose_num, which takes two positive numbers $x$ and $y$ and returns the largest even integer that falls into the interval $[x,y]$. We first use the “How to solve:” as a prompt, then break down the overall problem into multiple steps and describe the programming ideas in natural language. This includes explaining branching and looping structures to guide the model’s understanding of the task.
While these previous studies demonstrate the potential of CoTs in improving the performance of LLMs for code generation, they possess certain limitations as the current methods for CoT generation heavily rely on manual writing of CoTs or the utilization of LLMs [16, 17], leading to high costs. These limitations motivate us to investigate the following two main questions. (1) Can $\ell$LMs independently generate high-quality CoTs to guide code generation, and (2) can $\ell$LMs benefit from generated CoTs? Here, “independently” means no model training or model parameter updating.
Empirical observations. To address the first question, we conduct empirical studies on the CoT generation capabilities of 11 different $\ell$LMs and two LLMs. We adopt a few-shot approach, which provides $\ell$LMs with a set of examples to generate the corresponding CoT. Our finding shows that most $\ell$LMs with parameter scales ranging from 0.3 to 7 billion, unfortunately, did not demonstrate the ability to generate high-quality CoTs independently (cf. Section 5.1 for details). To address the second question, we compare the performance of $\ell$LMs in code generation with and without CoTs. Our findings suggest that all $\ell$LMs obtain performance improvement with the CoTs. As an example, the performance of the CodeT5 + 6B model on the HumanEval-plus dataset [4] can be improved from 26.83% to 43.90% with the CoTs generated by our methods (cf. Section 5.3 for details).
Figure 1 provides a motivation example, where CodeGen [18] is used as a case study. We evaluate its performance by considering varying parameter sizes 350M, 2B, and 6B. Without CoT, these models did not generate the correct code (cf. Fig 1(a)). However, with CoT, we decompose user requirements into three intermediate steps. In the first step, we initialize a variable max_even as -1; in the second step, we define the details of loop conditions and judgment conditions; in the third step, we return the value. As such, we can effectively instruct the models on the necessary actions at each step, and they eventually generate semantically correct code (though variables are with different names).
@@
Evaluation. We conduct a comprehensive evaluation of the quality of the CoTs generated by COTTON on the HumanEval benchmark [24]. To ensure the generalizability of COTTON, we further collected a new code generation dataset OpenEval, and evaluated COTTON on the OpenEval benchmark as well. Specifically, we selected nine other commonly used models as base models and compared the results with the same training process. The quality of CoTs generated by COTTON is superior to others in both automated and human evaluation metrics. Furthermore, we evaluate the performance on $\ell$LMs when adopting the generated CoTs on code generation benchmarks (such as HumanEval, HumanEval-plus, and OpenEval). The results show that, for various $\ell$LMs, the CoTs generated by COTTON achieve higher performance gains than those generated by LLMs such as ChatGLM (130B), and are competitive with those generated by gpt-3.5-turbo (175B).
@@ 2.1 Code Generation Task Formulation
In our study, we treat the existing neural code generation model $M_{code}$ as a black box. We intent to train a model $M_{cot}$ to generate CoTs, which can be used to guide code generation and further improve the performance of this task.
@@ 2.2 Code Language Model CodeLlama
where $SiLU$ represents the activation function, which is defined as the element-wise product of the Sigmoid function and the input. This activation function introduces non-linearity to the network.
@@ 3.1 Data Collection
    Doc Filtering. We utilize the DocChecker  [^7] https://github.com/FSoft-AI4Code/DocChecker to determine the consistency between the documentation and the code. We then remove code snippets with inconsistent documentation.
@@
    Quality Checker. This agent assesses the educational value of the data and removes low-quality items, ensuring that the dataset comprises high-quality code snippets. The specific prompt is illustrated as follows.
    Give you a code snippet, determine its educational value for a student whose goal is to learn basic coding concepts. If it has educational value, return only “Yes”, else return ”No”.
@@
    CoT Generator. We first transform code snippets and functional comments into a standardized prompt and solution format similar to HumanEval [24]. This agent employs a one-shot approach, providing an example to aid the agent in learning the desired output style and generating CoT based on user input. Importantly, we intentionally do not disclose the specific implementation details of the code to the agent, which encourages the agent to generate diverse CoTs, as code implementation can vary widely. The specific prompt is illustrated as follows.
    ### Given a piece of code, output the corresponding implementation idea. ### Example: Input: from typing import List def below_zero(operations: List[int]) ->bool: ””” You’re given a list of deposit and withdrawal operations on a bank account that starts with zero balance. Your task is to detect if at any point the balance of account falls below zero, and at that point function should return True. Otherwise it should return False. ””” Output: How to solve: Step 1. Initialize account balance as 0. Step 2. Iterate through operations. -add value to account balance. -If account balance <0, return True. Step 3. Return False. ### Input: $[X]$## Output: $[Y]$
@@
    Consistency Checker. This agent examines the consistency between the CoT instructions generated by Agent 2 and the code snippets. It removes code snippets with inconsistencies, ensuring that the CoT instructions accurately reflect the code’s behavior. The specific prompt is illustrated as follows.
    Given a piece of code and a chain of thought, determine whether they express exactly the same functional semantics. If consistent, return only ”Yes”, else return ”No”.
@@ 3.2 Model Training
To enable the training of $\ell$LMs with limited resources, researchers have explored parameter-efficient fine-tuning methods since full parameter fine-tuning is impractical in this scenario. As demonstrated in the previous study [33], these methods have been proven to achieve high performance by providing sufficient instructions to the language models. In contrast to continuous-based soft prompt methods [34, 35], our approach employs a set of discrete tokens as prompts, which are meaningful and easily interpretable. For our CoT generation task, we design a template that incorporates task-specific instructions, which is illustrated as follows.
### Given a piece of code, output the corresponding implementation idea. ### Input: $[X]$## Output: $[Y]$
@@ 3.3 Model Inference
In the inference phase, the model trained by COTTON can be efficiently deployed on a single consumer graphics card to generate CoT. Note that COTTON is a stand-along tool and its deployment does not require the use of LLM. To accelerate the decoding process, COTTON utilizes the Greedy Search algorithm, which, in a nutshell, selects the token with the highest probability at each decoding step, resulting in a more deterministic output during inference.
@@ 4.1.1 Code Generation
OpenEval. To ensure fairness and generalizability, we collect a new code generation dataset OpenEval. This dataset includes 178 problems selected from the competition-level code translation dataset AVATAR [36]. For each problem, we designed five test cases in a manual way that can effectively evaluate the quality of the generated code and minimize bias and leakage.
@@ 4.4 Implementation Details and Running Platform
Learning Rate 5e-5 Training batch size 16
Beam size 10 Validation batch size 16
@@
Our implementation is based on PyTorch 1.8, and the experiments are conducted on a machine with an Intel(R) Xeon(R) Silver 4210 CPU and the GeForce RTX 3090 GPU.
@@ 5.1 RQ1: Can $\ell$LMs generate high-quality CoTs independently?
To investigate whether $\ell$LMs can independently generate high-quality CoTs for code generation, we conduct experiments based on few-shot prompt learning techniques as described in Section 3.1. We evaluate the performance of $\ell$LMs with model sizes ranging from 0.35B to 7B, the base model CodeLlama, and LLMs (such as InternLM 123B [48] and ChatGLM 130B).
Table III presents the performance of $\ell$LMs and LLMs across all evaluation metrics on the HumanEval-CoT and OpenEval-CoT datasets. In terms of lexical similarity, $\ell$LMs generally performed worse than LLMs. For example, using the METEOR metric, InternLM 123B and ChatGLM 130B achieved scores around 0.32 on the HumanEval-CoT dataset and around 0.35 on the OpenEval-CoT dataset. In contrast, the majority of $\ell$LMs scored below 0.3 on both datasets. In terms of semantic perspective, most $\ell$LMs are difficult to generate high-quality CoTs, while LLMs demonstrated better performance. For example, based on the Consistency metric, InternLM 123B and ChatGLM 130B achieved scores above 0.86 on the HumanEval-CoT dataset and above 0.76 on the OpenEval-CoT dataset. In contrast, the majority of $\ell$LMs scored below 0.6 on both datasets.
@@
The majority of $\ell$LMs face challenges in generating high-quality CoTs for guiding code generation independently.
TABLE IV: Performance comparison of different base models to generate CoTs on the HumanEval-CoT and OpenEval-CoT datasets
@@ 5.2 RQ2: Can COTTON generate higher-quality CoTs
Automatic Evaluation. We first conduct an evaluation of different base models on the HumanEval-CoT and OpenEval-CoT datasets. We consider different evaluation metrics, including BLEU-1, BLEU-2, BLEU-3, BLEU-4, Meteor, Rouge-L, and Consistency. The results are shown in Table IV.
For the HumanEval-CoT dataset, COTTON consistently outperformed the other base models across various evaluation metrics. This indicates that COTTON generates CoTs with higher lexical similarity compared to the base models. For example, based on the METEOR metric, COTTON achieved a score of 0.38, while the other base models scored between 0.27 and 0.37. In terms of semantic similarity, COTTON also outperformed the other base models on the HumanEval-CoT dataset. With a score of 0.93, COTTON demonstrated a higher level of semantic similarity to the actual code compared to the scores ranging from 0.29 to 0.92 achieved by the other base models.
Importantly, COTTON even outperforms larger LLMs (i.e., InternLM 123B and ChatGLM 130B), indicating its effectiveness in generating high-quality CoTs. We can find similar trends on the OpenEval-CoT dataset in Table IV, where COTTON achieves better performance compared to the other base models.
@@
Human Evaluation. Automatic evaluation metrics can provide valuable insights into the quality of generated CoTs, However, these metrics primarily focus on overlap or semantic similarity with the CoTs generated by the Teacher Model, but they may not truly reflect whether the generated CoT is inspiring or provides meaningful guidance to developers.
To further assess the quality of the CoTs generated by CodeGeeX2 and our proposed approach COTTON, we conducted a human study. This study involved human evaluators who provided subjective evaluation and insights for the practical usability and effectiveness of the generated CoTs. By involving human evaluators, we aimed to obtain a more comprehensive assessment of the CoTs and gain a deeper understanding of their practical implications.
@@
In our human study, we adopted the methodology used in previous studies [56, 57]. The quality of the generated CoTs was evaluated based on three aspects, as illustrated in Figure 4:
@@
For the human evaluation, we recruited six master students who had over six years of project development experience and were familiar with Python programming language. We selected all samples from the HumanEval-CoT dataset (164 samples) and the OpenEval-CoT dataset (178 samples) as the evaluation subjects. For each sample, we collected the ground truth CoT and two generated CoTs, one from CodeGeeX2 and one from our proposed COTTON.
To ensure a thorough evaluation, we divided all the samples into three groups, with each group containing 114 samples. Each group was evaluated by two master students (assessors) in terms of naturalness and educational value. The score for each aspect ranged from 0 to 4, with higher scores indicating higher quality. Each generated CoT was evaluated by two different master students, and we took the average score value.
A sample questionnaire is shown in Figure 4. To guarantee human study quality, the generated CoTs were presented in a random order, ensuring that the assessors had no knowledge of which approach generated the CoT. Moreover, the assessors were allowed to use the internet to look up any related concepts they were unfamiliar with. Finally, we limited each master student to evaluate only 20 samples in half a day. This was done to prevent fatigue and maintain a high level of concentration during the evaluation process.
TABLE V: The average score and standard deviation (in parentheses) of human study
Dataset Aspect CodeGeeX2 COTTON
HumanEval-CoT Similarity 2.762 (0.918) 3.038 (0.743)
Naturalness 3.375 (0.592) 3.442 (0.713)
Educational Value 2.973 (0.812) 3.102 (0.779)
OpenEval-CoT Similarity 2.976 (0.615) 3.128 (0.734)
Naturalness 3.567 (0.789) 3.691 (0.657)
Educational Value 3.002 (0.876) 3.209 (0.891)
Our human study results are presented in Table V, where the average score and standard deviation of all the samples evaluated by the master students are shown. The results show that our proposed approach COTTON outperforms CodeGeeX2 in all three perspectives: similarity, naturalness, and educational value. These findings provide further evidence of the effectiveness of COTTON in generating high-quality CoTs for guiding code generation. The higher average score values obtained by COTTON indicate that it is capable of generating CoTs that are more grammatically correct, fluent, and valuable in terms of providing guidance and inspiration to developers. These results highlight the competitiveness of our proposed approach in generating CoTs that are not only grammatically sound but also valuable from an educational perspective.
COTTON consistently outperforms the state-of-the-art base models in generating CoTs in terms of lexical and semantic similarity, as well as its ability to provide valuable guidance and inspiration to developers.
TABLE VI: The performance of code generation models with or without the CoTs generated by different methods on the HumanEval(HE), HumanEval-plus(HE-p), and OpenEval(OE) datasets

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable potential in code generation. The integration of Chain of Thought (CoT) reasoning can further boost their performance. However, current CoT methods often require manual writing or LLMs with over 100 billion parameters to generate, impeding their applicability in resource-constrained scenarios. In this study, we investigate lightweight Language Models ($\ell$LMs), which are defined to have fewer than 10 billion parameters. Empirically, we find that most $\ell$LMs cannot generate high-quality CoTs when prompted by the few-shot method, but can take advantage of high-quality CoTs generated elsewhere to improve their performance in code generation. Based on these findings, we design a novel approach COTTON which can leverage $\ell$LMs to automatically generate CoTs for code generation. We synthesize new datasets and conduct extensive experiments on various benchmarks. The results show that the CoTs generated by COTTON outperform the baselines in terms of automated and human evaluation metrics. In particular, the CoTs generated by COTTON boost various $\ell$LMs to achieve higher performance gains than those generated by LLMs such as ChatGLM (130B), and are competitive with those generated by Gemini and gpt-3.5-turbo. The results also reveal that COTTON not only improves the performance of $\ell$LMs, but also enhances the performance of LLMs. Our study showcases the potential of $\ell$LMs in software engineering applications.
@@ 1 Introduction
Neural code generation, which can automatically generate programs from natural language requirements based on deep learning, has become a promising approach to meet the challenges of the ever-increasing complexity of software and alleviate the burden on programmers [1, 2]. Recently, large language models (LLMs), such as GPT4 [3], have demonstrated impressive performance in code generation tasks [4]. The state-of-the-art LLMs normally have over 100 billion parameters, making even their deployment highly non-trivial. These LLMs pose challenges in terms of time, computational, and financial costs when applied to code generation, rendering them impractical for most individual users, or in resource-constrained scenarios, such as restricted access to LLM APIs or constrained GPU availability. [5, 6]. For software engineering applications, it is imperative to develop lightweight language-model-based techniques that are more friendly for users (e.g., individual end users). Fu et al. [7] defined models with parameters greater than 100B as large models and those with parameters less than 10B as small models. Admittedly, the precise definition of large and small models is debatable and may evolve with the advance of technology. In this study, we define (pre-trained) language models (LMs) with less than 10 billion parameters as lightweight Language Models ($\ell$LM), the rationale of which is that these models can be deployed on a single user graphics card (e.g., RTX 3090 or RTX 4090) based on the current technology. The general aim is to develop techniques to tackle software engineering challenges based on $\ell$LMs but with competitive performance as state-of-the-art LLMs, which would enable efficient, yet more accessible, software engineering applications.
@@
The previous research has primarily focused on investigating the impact of CoT on LLMs, leaving questions regarding whether $\ell$LMs can also benefit from the guidance of CoT. In Figure 1, we present a motivation example to demonstrate the potential of CoT for $\ell$LMs in code generation. Specifically, in Figure1(a), the programming task is choose_num, which takes two positive numbers $x$ and $y$ and returns the largest even integer that falls into the interval $[x,y]$. The example highlights that the original prompts for the $\ell$LMs (CodeGen-350M, CodeGen-2B, and CodeGen-6B) fail to generate correct code solutions. However, by leveraging CoT in Figure1(b), we modify the original prompt by using ”How to solve:” and breaking down the problem into multiple steps, where the natural language explanations guide the model’s understanding of the task, including instructions on branching and looping structures. With the new CoT, these $\ell$LMs can generate correct code solutions.
Furthermore, previous studies possess certain limitations as the current methods for CoT generation heavily rely on manual writing of CoTs or the utilization of LLMs [16, 17], leading to high costs. These limitations motivate us to investigate the following two main questions. (1) Can $\ell$LMs independently generate high-quality CoTs to guide code generation, and (2) can $\ell$LMs benefit from generated CoTs? Here, “independently” means no model training or model parameter updating.
Empirical observations. To address the first question, we conduct empirical studies on the CoT generation capabilities of 11 different $\ell$LMs and two LLMs. We adopt a zero-shot approach [18] and some few-shot approaches (such as Self-planning [13], SCoT [14], and the self-cot we propose), which provide $\ell$LMs with a set of examples to generate the corresponding CoT. Our finding shows that most $\ell$LMs with parameter scales ranging from 0.3 to 7 billion, unfortunately, did not demonstrate the ability to generate high-quality CoTs independently (cf. Section 5.1 for details). To address the second question, we compare the performance of $\ell$LMs in code generation with and without CoTs. Our findings suggest that all $\ell$LMs obtain performance improvement with the CoTs. As an example, the performance of the CodeT5 + 6B model on the HumanEval-plus dataset [4] can be improved from 26.83% to 43.90% with the CoTs generated by our methods (cf. Section 5.3 for details).
Figure 1 provides a motivation example, where CodeGen [19] is used as a case study. We evaluate its performance by considering varying parameter sizes 350M, 2B, and 6B. Without CoT, these models did not generate the correct code (cf. Fig 1(a)). However, with CoT, we decompose user requirements into three intermediate steps. In the first step, we initialize a variable max_even as -1; in the second step, we define the details of loop conditions and judgment conditions; in the third step, we return the value. As such, we can effectively instruct the models on the necessary actions at each step, and they eventually generate semantically correct code (though these variables have different names).
@@
Evaluation. We conduct a comprehensive evaluation of the quality of the CoTs generated by COTTON on the HumanEval benchmark [25]. To ensure the generalizability of COTTON, we further collected a new code generation dataset OpenEval, and evaluated COTTON on the OpenEval benchmark as well. Specifically, we selected nine other commonly used models as base models and compared the results with the same training process. The quality of CoTs generated by COTTON is superior to others in both automated and human evaluation metrics. Furthermore, we evaluate the performance on $\ell$LMs when adopting the generated CoTs on code generation benchmarks (such as HumanEval, HumanEval-plus, and OpenEval). The results show that, for various $\ell$LMs, the CoTs generated by COTTON achieve higher performance gains than those generated by LLMs such as ChatGLM (130B) [26], and are competitive with those generated by Gemini and gpt-3.5-turbo.
@@
In addition, to further evaluate the capabilities of COTTON, we conducted experiments to assess its impact on LLMs (such as gpt-3.5-turbo). The results show that gpt-3.5-turbo could show a significant performance improvement in code generation when guided by the CoTs generated by COTTON where the performance even exceeds that of the GPT-4 zero-shot scenario on the HumanEval dataset.
Finally, to demonstrate the effectiveness of COTTON compared to fine-tuning, we utilize the StarCoder-series models as examples. The results indicate that the combination of StarCoder-7B with COTTON has already exceeded the performance of StarCoder-16B in zero-shot scenarios and can even achieve comparable results to a fine-tuned StarCoder-16B model. The efficacy of COTTON in enhancing performance across multiple models without the necessity of fine-tuning individual model is noteworthy, as it not only saves time and computation resources for model construction but also opens an inviting avenue to avoid traditional fine-tuning in adapting language models.
@@ 2.1 Code Generation Task Formulation
In our study, we treat the existing neural code generation model $M_{code}$ as a black box. We intend to train a model $M_{cot}$ to generate CoTs, which can be used to guide code generation and further improve the performance of this task.
@@ 2.2 Code Language Model CodeLlama
where $SiLU$ represents the activation function, defined as the element-wise product of the Sigmoid function and the input. This activation function introduces non-linearity to the network.
@@ 3.1 Data Collection
    Doc Filtering. We utilize DocChecker [34] to determine the consistency between the documentation and the code, which can maintain an accurate alignment between the comments and the code effectively. We then remove code snippets with inconsistent documentation.
@@
    Quality Checker. This agent assesses the educational value of the data and removes low-quality items, ensuring that the dataset comprises high-quality code snippets.
@@
    CoT Generator. We first transform code snippets and functional comments into a standardized prompt and solution format similar to HumanEval [25]. This agent employs a one-shot approach, providing an example to aid the agent in learning the desired output style and generating CoT based on user input. Importantly, we intentionally do not disclose the specific implementation details of the code to the agent, which encourages the agent to generate diverse CoTs, as code implementation can vary widely.
@@
    Consistency Checker. This agent examines the consistency between the CoT instructions generated by Agent 2 and the code snippets. It removes code snippets with inconsistencies, ensuring that the CoT instructions accurately reflect the code’s behavior.
Quality Checker Give you a code snippet, determine its educational value for a student whose goal is to learn basic coding concepts. If it has educational value, return only “Yes”, else return ”No”.
Consistency Checker Given a piece of code and a chain of thought, determine whether they express exactly the same functional semantics. If consistent, return only ”Yes”, else return ”No”.
CoT Generator ### Given a piece of code, output the corresponding implementation idea. ### Example: Input: from typing import List def below_zero(operations: List[int]) ->bool: ””” You’re given a list of deposit and withdrawal operations on a bank account that starts with zero balance. Your task is to detect if at any point the balance of account falls below zero, and at that point function should return True. Otherwise it should return False. ””” Output: How to solve: Step 1. Initialize account balance as 0. Step 2. Iterate through operations. -add value to account balance. -If account balance <0, return True. Step 3. Return False. ### Input: $[X]$## Output: $[Y]$
@@ 3.2 Model Training
To enable the training of $\ell$LMs with limited resources, researchers have explored parameter-efficient fine-tuning methods since full parameter fine-tuning is impractical in this scenario. As demonstrated in the previous study [36], these methods have been proven to achieve high performance by providing sufficient instructions to the language models. In contrast to continuous-based soft prompt methods [37, 38] , our approach employs a set of discrete tokens as instruction prompts, which are meaningful and easily interpretable. For our CoT generation task, we design a template that incorporates task-specific instructions, which is illustrated as follows.
Instruction Template ### Given a piece of code, output the corresponding implementation idea. ### Input: $[X]$## Output: $[Y]$
@@ 3.3 Model Inference
In the inference phase, the model trained by COTTON can be efficiently deployed on a single consumer graphics card to generate CoT. Note that COTTON is a stand-alone tool and its deployment does not require the use of LLM. To accelerate the decoding process, COTTON utilizes the Greedy Search algorithm, which, in a nutshell, selects the token with the highest probability at each decoding step, resulting in a more deterministic output during inference.
@@ 4.1.1 Code Generation
OpenEval. To ensure fairness and generalizability, we collect a new code generation dataset OpenEval. This dataset includes 178 problems selected from the competition-level code translation dataset AVATAR [39]. For each problem, we designed additional test cases in a manual way that can effectively evaluate the quality of the generated code and minimize bias and leakage. Particularly, we hired two software engineers with 2$\sim$3 years of development experience each to construct 5 test cases for each code segment to ensure the diversity of test cases.
@@ 4.4 Implementation Details and Running Platform
Learning Rate 1e-4 Training batch size 1
@@
Epoch 20 Early Stop 5
Our implementation is based on PyTorch 1.8, and the experiments are conducted on a machine with an Intel(R) Xeon(R) Silver 4210 CPU and the GeForce RTX 3090 GPU. It took about 6 hours to complete the model training of COTTON.
@@ 5 Experimental Result Analysis
Gemini Not-Available 52.50 43.56 36.43 30.92 32.14 53.28 91.46
gpt-4 Not-Available 60.14 50.49 43.03 37.24 34.55 54.32 96.95
Teacher(gpt-3.5-turbo) Not-Available - - - - - - 93.29
@@
Gemini Not-Available 57.49 49.04 42.43 37.25 34.13 54.88 81.46
gpt-4 Not-Available 62.54 54.34 47.88 42.81 36.84 58.02 87.64
Teacher(gpt-3.5-turbo) Not-Available - - - - - - 89.33
@@ 5.1 RQ1: Can $\ell$LMs generate high-quality CoTs independently?
To investigate whether $\ell$LMs can independently generate high-quality CoTs for code generation, we conduct experiments based on few-shot prompt learning techniques as described in Section 3.1. We evaluate the performance of $\ell$LMs with model sizes ranging from 0.35B to 7B, the base model CodeLlama, and several representative LLMs (such as InternLM 123B [51], ChatGLM 130B [26], Gemini [52], gpt-3.5-turbo, and gpt-4 [53]). Among these LLMs, Gemini, gpt-3.5-turbo, and gpt-4 are currently considered to be the most promising LLMs in code generation tasks and have shown promising performance.
Table III presents the performance of $\ell$LMs and LLMs across all evaluation metrics on the HumanEval-CoT and OpenEval-CoT datasets. In terms of lexical similarity, $\ell$LMs generally performed worse than LLMs. For example, using the METEOR metric, LLMs like gpt-4 can achieve scores around 0.35 on the HumanEval-CoT dataset and around 0.37 on the OpenEval-CoT dataset. In contrast, the majority of $\ell$LMs scored below 0.3 on both datasets. In terms of semantic perspective, most $\ell$LMs are difficult to generate high-quality CoTs, while LLMs demonstrated better performance. For example, based on the Consistency metric, gpt-4 can achieve scores above 0.96 on the HumanEval-CoT dataset and above 0.87 on the OpenEval-CoT dataset. In contrast, the majority of $\ell$LMs scored below 0.6 in both data sets. In addition, in terms of the Consistency metric, gpt-3.5-turbo and gpt-4 achieved the best results on the OpenEval and HumanEval datasets, respectively, yet the cost of calling gpt-4 is 20 times higher than that of gpt-3.5-turbo. Therefore, we choose gpt-3.5-turbo as the Teacher Model in our study.
@@
Summary of RQ1 The majority of $\ell$LMs face challenges in generating high-quality CoTs for guiding code generation independently.
TABLE IV: Performance comparison of different base models to generate CoTs on the HumanEval-CoT and OpenEval-CoT datasets (Here COTTON is based on CodeLlama-7b).
@@ 5.2 RQ2: Can COTTON generate higher-quality CoTs
5.2.1 Automatic Evaluation.
We first evaluate different base models on the HumanEval-CoT and OpenEval-CoT datasets. We consider different evaluation metrics, including BLEU-1, BLEU-2, BLEU-3, BLEU-4, Meteor, Rouge-L, and Consistency. The results are shown in Table IV.
For the HumanEval-CoT dataset, COTTON consistently outperformed the other base models across various evaluation metrics. This indicates that COTTON generates CoTs with higher lexical similarity compared to the base models. For example, based on the METEOR metric, COTTON achieved a score of 0.38, while the other base models scored between 0.27 and 0.37. In terms of semantic similarity, COTTON also outperformed the other base models on the HumanEval-CoT dataset. With a score of 0.93, COTTON demonstrated a higher level of semantic similarity to the actual code compared to the scores ranging from 0.29 to 0.92 achieved by the other base models. Importantly, COTTON even outperforms larger LLMs (i.e., InternLM 123B, ChatGLM 130B, and Gemini), indicating its effectiveness in generating high-quality CoTs. We can find similar trends on the OpenEval-CoT dataset in Table IV, where COTTON achieves better performance compared to the other base models.
@@
5.2.2 Human Evaluation.
Although automatic evaluation metrics can offer valuable insights into the quality of generated CoTs, these metrics mainly concentrate on overlap or semantic similarity with the CoTs generated by the Teacher Model, possibly overlooking whether the generated CoT is indeed inspiring or offers meaningful guidance to developers. Furthermore, the enhancements in Table IV brought about by COTTON over CodeGeeX2 and LLama2 may not be immediately apparent.
To further evaluate the quality of the generated CoTs, we conducted a human study using four groups of CoTs generated by different base models (CodeGeeX2 and LLama2) that performed similarly based on the results in Table III and Table IV. By involving human evaluators, we aimed to obtain a more comprehensive assessment of the CoTs and gain a deeper understanding of their practical implications.
@@
In our human study, we adopted the methodology used in previous studies [61, 62, 63]. The quality of the generated CoTs was evaluated based on three aspects, as illustrated in Figure 4:
@@
TABLE V: The average score and standard deviation (in parentheses) of human study.
Aspect LLama2 CodeGeeX2 COTTON
HE Similarity 2.847 (0.875) 2.762 (0.918) 3.038 (0.743)
Naturalness 3.418 (0.842) 3.375 (0.592) 3.442 (0.713)
Educational Value 2.985 (0.788) 2.973 (0.812) 3.102 (0.779)
OE Similarity 3.024 (0.645) 2.976 (0.615) 3.128 (0.734)
Naturalness 3.521 (0.746) 3.567 (0.789) 3.691 (0.657)
Educational Value 2.955 (0.858) 3.002 (0.876) 3.209 (0.891)
TABLE VI: The p-values between three aspects of the human study, and all the p-values are substantially smaller than 0.005.
Model Similarity Similarity Educational Value
HE Llama2 2.14e-5 1.85e-5 4.23e-5
CodeGeeX2 5.68e-5 3.62e-5 1.45e-5
OE Llama2 1.53e-4 4.28e-4 8.85e-4
CodeGeeX2 1.68e-4 4.29e-4 3.67e-4
TABLE VII: The performance of code generation models with or without the CoTs generated by different self sot prompt methods on the HumanEval(HE), HumanEval-plus(HE-p), and OpenEval(OE) datasets